OpenStudy (anonymous):

What is the solution to the equation 2k - 10 ------- =8 -7 k= ?

TheSmartOne (thesmartone):

why, hello there!

TheSmartOne (thesmartone):

Is your question \(\sf \Large \frac{2k-10}{-7} = 8\)

OpenStudy (error1603):

k-9 = 0

OpenStudy (error1603):

Solve

OpenStudy (error1603):

Add 9 to both sides of the equation

OpenStudy (error1603):

k = 9

TheSmartOne (thesmartone):

Error, seems like you have an error.

OpenStudy (alekos):

ha ha that's what i reckon too

OpenStudy (error1603):

What is it?

TheSmartOne (thesmartone):

#1, you're supposed to let the asker participate in the discussion :)

OpenStudy (anonymous):

this is the question

TheSmartOne (thesmartone):

ok, lets multiply -7 on both sides \(\sf \Large \frac{2k-10}{-7}\times -7= 8\times -7\) Simplify it :) What do you get?

OpenStudy (anonymous):

huh i dont what u just typed

OpenStudy (anonymous):

someone please help

TheSmartOne (thesmartone):

if you don't cooperate you won't get help

TheSmartOne (thesmartone):

tell us what you don't get :)